Ricks paint store sells a special shade of pink paint (Ricky's Blush) which can be made from red paint and white paint. The amount of red paint should be three times the amount of white paint. How much of each kind of paint is needed to make 264 gallons of Ricky's Blush?
Answer by Maths68(1474)

Amount of White paint is used = x
The amount of red paint should be three times the amount of white paint
Amount of Red paint is used = 3x
Amount of Riky's Blush = 264


White paint + Red paint = Riky's Blush
x+3x=264
4x=264
4x/4=264/4
x=66



Amount of White paint is used = x =66 gallons
The amount of red paint should be three times the amount of white paint
Amount of Red paint is used = 3x = 3*66 = 198 gallons